what is the LR when 10mol iron sulfide reacts 200g oxygen FeS2+o2


1
Expert's answer
2021-02-24T05:31:26-0500

The chemical reaction is as follows:

4 FeS2 + 11 O2 = 2 Fe2O3 + 8 SO2

From here:

n(FeS2) = 10 mol / 4 = 2.5 mol

n(O2) = m(O2) / (11 × Mr[O2]) = 200 g / (11 × 32) = 0.57 mol

Since, the number of moles of O2 is lower than the number of moles of FeS2, oxygen is a limiting reactant.


Answer: Oxygen (O2)